# TraveLER: A Multi-LMM Agent Framework for Video Question-Answering

### v1

@@ 4.1 Datasets
We use the following datasets: (1) NExT-QA (Xiao et al., 2021) is a dataset that tests causal action reasoning and temporal understanding, with questions categorized into a Temporal, Causal, or Descriptive type. NExT-QA requires going beyond simple recognition of objects and actions to answer the questions correctly. Following the trend of works before us, we evaluate our method on the 5,000 questions in the NExT-QA validation set. (2) Perception Test (Pătrăucean et al., 2023) is a dataset that focuses on skills such as memory, abstraction, physics, and semantics and is intended to be approached in a few-shot or zero-shot manner. The dataset consists of 11.6k real-world videos and 38K multiple-choice QA questions. (3) STAR (Wu et al., 2021) is a dataset that tests reasoning in real-world video situations. It consists of 22K video clips, with 60K situated reasoning questions. Questions are broadly divided into 4 main categories: interaction, sequence, prediction, and feasibility.
@@ 4.2 Baselines
In our experiments, we compare our method to recent state-of-the-art zero-shot (ZS) methods, such as LLoVi (Zhang et al., 2023a), ProViQ (Choudhury et al., 2023), and other methods that are not necessarily ZS, such as SeViLA (Yu et al., 2023), and MC-ViT (Balažević et al., 2024). We note that SeViLA uses fine-tuned components on QV-Highlights (Lei et al., 2021a) in their model, while the MC-ViT model is fine-tuned on NExT-QA for Perception Test.
@@ 4.3 Results
Our results are shown in Table 1, Table 3, and Table 3. First, we see that on NExT-QA, our method outperforms LLoVi using GPT-4, which is the current state-of-the-art that uniformly captions frames across the entire video. Interestingly, our method demonstrates superior performance in comparison to LLoVi despite viewing 50% fewer frames on average, highlighting the effectiveness of our approach. Second, we also outperform SeViLA by +4.6%, although SeViLA uses a keyframe selector that is fine-tuned on a video moment retrieval/grounding task. Lastly, we see comparable performance to BLIP-2voting (60.5% vs 61.3%) on the Temporal split, even though BLIP-2voting views every single frame.
@@
For Perception Test and STAR, we use GPT-3.5 because it is much cheaper than GPT-4, but results are likely to be improved even further with GPT-4. Nevertheless, we achieve higher accuracy than LongViViT on Perception Test by +4.5% and MC-ViT by +2.1%, although it was fine-tuned on NExT-QA. We surpass on STAR both the best zero-shot approach by +2.7% and the best fine-tuned result by +0.3%.
@@
Finally, please refer to Section D for visualizations. For example, in Figure 8, we see that question-answering is able to extract more relevant details in comparison to simple captioning. This may explain why our method significantly outperforms the descriptive split of NeXT-QA. For Perception Test, the questions require a high-level understanding of the video and fine-grained details. In Figure 7, we see that our method is able to reason about vague references and correctly identify relevant objects through question-answering.
@@
Table 4: Results of different ablations on our framework. We report (a) replacing different LLMs and LMMs, (b) selecting different numbers of frames to view in the Retriever, and (c) changing the number of questions asked in the Extractor.
((a)) Replacing Diff. Agents
@@
4.4 Ablations
We perform comprehensive ablations using 1000 randomly selected questions from the NExT-QA training set to understand the impact of each component. Unless otherwise specified, we use GPT-3.5 as the LLM and LLaVA-1.6 as the LMM for all agents.
@@ 4.4 Ablations
Substituting different agents. In our method, we use LLMs and LMMs as the main agents. To understand how the choice of the LLM and LMM affects our framework’s performance, we swap different LLMs and LMMs into our framework (see Table 4(a)). First, we examine the performance of our framework using different LLMs while fixing the LMM. For this, we use LLaVA-1.6 as our LMM, and we find that GPT-4 performs the best in our framework by a significant margin of +5.4% compared to GPT-3.5. We hypothesize that stronger LLMs perform better in our framework due to how the LLM components are required to perform advanced reasoning tasks. Second, we measure the performance of our framework using different LMMs for $\text{LLM}_{\text{extractor}}$ while fixing the LLM to be GPT-3.5. We find that LLaVA-1.6 performs best, GPT-4V is slightly worse (-0.9%), and BLIP-2 is significantly worse (-7.7%).
Ablating the Planner. The Planner module outputs a plan, a list of instructions that guides the behavior of all other modules. In order to evaluate the effectiveness of this module, we remove it from our framework. We find that removing the Planner leads to a drop in accuracy of -2.3%. This may be due to several reasons. First, the Planner provides many temporal cues that guide the Retriever module’s search, such as “go to the middle of the video”, and without these cues, the Retriever is not as good at selecting the next timestamp. Moreover, the Planner also helps the Evaluator better decide when to stop since in our iterative approach, the Evaluator uses the plan to determine when to stop.
Ablating the Retriever. The Retriever module determines the next timestamps to view, which helps focus our traversal and information collection. To evaluate the effectiveness of the Retriever, we uniformly sample frames from the video at 2-second intervals, similar to other methods, such as LLoVi (Zhang et al., 2023a). We find that this leads to a reduction in overall accuracy by -3.5%. Our results suggest that compared to uniform sampling, the Retriever allows us to capture frames that might have otherwise been skipped. This also indicates that our Retriever selects fewer unimportant frames that might mislead the model.
Ablating the Extractor. Question asking is a key component of our model as it allows us to capture more fine-grained and question-relevant information in comparison to simple caption generation, which produces a generic description. As such, we ablate the Extractor by only allowing the LMM to caption frames. We find that this decreases performance by -2.2%, suggesting that the ability to ask specific questions about a frame is important. We notice that many generated captions capture the main idea of visual information in the frame, but often miss fine-grained details that are useful for answering the question.
@@
Retriever window size. To evaluate the impact of window size, which is the number of frames the Retriever extracts centered around the selected frame, we experiment with multiple-frame retrieval. This allows to capture better actions that occur quickly or require more context to understand. As shown in Table 4(b), we find that choosing 5 frames yields the best results and a +1.4% increase when compared to selecting a single frame, but viewing more than 5 decreases performance. This suggests that the ability to retrieve multiple frames is beneficial in allowing the model to better capture relevant information, but retrieving too many frames can lead to too much information, resulting in performance drops.
@@
The number of questions. Question answering allows us to extract more specific details from our visual inputs. However, we noticed that too many questions can yield irrelevant questions and false positives. As such, we experiment with modifying the number of questions asked for each frame by our extractor (see Table 4(c)). We record results for a 5-question, 3-question, and 1-question maximum. Note that 0-questions asked is equivalent to only allowing captions, which is discussed in the Extractor ablation. From our results, we find that a 3-question limit yields the best results compared to asking 1 or 5 questions (+2.0%/+0.8%). This suggests that asking questions helps in extracting relevant information, but too many questions can lead to false positives or too much irrelevant information.
@@
Memory bank initialization and formatting. Our memory bank $M$ is a critical part of our framework since it stores information that all modules rely on to evaluate inputs and make decisions. First, we experiment with different initializations as $M$ must be initialized in the first iteration. We experiment with initialization of 1, 3, and 5 uniformly sampled captions. We find that using 5 evenly spaced frames yields the best results, possibly because it provides the model with a general overview of the video at the start before starting to collect more relevant information. Next, we experiment with changing the format of the memory bank, as there are many different ways to represent information. We try a markdown table instead of our dictionary format. Our results show that this performs worse by -2.9%, suggesting the effectiveness of our original dictionary formatting.
5 Discussion
Our TraveLER framework has demonstrated significant potential in utilizing an LMM image-based approach for VideoQA. We introduce a multi-LMM agent framework that travels along the video, collecting relevant information from keyframes through interactive question-asking. Our method creates a plan to “traverse” through the video, asking questions about individual frames to “locate” and store key information, and then “evaluate” if there is enough information to answer the question. Finally, if there is not enough information, our model is able to “replan” according to its collected knowledge. However, there are a few limitations to our work. Firstly, our framework depends on the strength of the LLM and LMM. We notice that false positives and incorrect statements from the LMM can impact performance. We also found that our method has a high runtime with slower LLMs since each iteration requires our LLM to generate significant amounts of text. We believe that with better and faster LLMs and LMMs in the future, these issues can be overcome. Finally, our research encourages future work on using large models for modular video approaches.
Acknowledgements
@@
We would like to thank Suzie Petryk, Chancharik Mitra, Alon Mendelson, and David Chan for helpful feedback and discussions. This project was supported in part by DoD, including PTG and/or LwLL programs, as well as BAIR’s industrial alliance programs.
@@ A.1 Additional Ablations
Memory Intialization. In order for the Planner to create effective plans, it is beneficial to initialize the memory bank properly. Memory initialization allows the Planner to have a high-level overview of the video, and create a corresponding plan on how to traverse the video given the initial frames. We perform three different initializations with 1, 3, and 5 frames and and display our results in Fig 3. We observe that initializing the memory bank with 5 frames (0, 0.25, 0.5, 0.75, 1 / beginning, quarter, middle, three-quarter, end) yields the best result. In contrast, we notice a decrease in accuracy by -0.6% when initializing with 3 frames (0, 0.5, 1 / beginning, middle, end) and a decrease by -2.5% when initializing with 1 frame (0.5 / middle).
Figure 4: Comparison of different LMM Response Length (75, 150, 300 max tokens). 150 tokens is optimal.
@@ A.2 Additional Results
We compare our method with other keyframe localization methods to further demonstrate the benefits of our TraveLER approach.
Method NExT-QA (Random Subset)
Tem. Cau. Des. Avg.
SeViLA - Localizer (Yu et al., 2023) 48.8 61.2 68.3 58.2
Moment-DETR (Lei et al., 2021a) 45.3 55.8 70.8 54.6
SigLIP (Zhai et al., 2023) 48.4 61.5 73.8 59.1
TraveLER - Planner & Retriever (ours) 50.9 62.7 72.4 60.3
@@ Appendix B Additional Implementation Details
To run our models on larger benchmarks, we use 8 NVIDIA RTX 6000 GPUs and split the dataset across multiple processes. In addition, we use the SGLang (Zheng et al., 2023) package, which provides a variety of performance optimizations for our LMMs and enables us to perform batched inference for models that do not natively support doing so. We serve our LMM on a single GPU and implement a queue that is shared across all runs. This allows individual runs to asynchronously call the LMM using an API request instead of creating a new instance of the LMM for each run. Typically, we have 4-5 processes sharing the same LMM at a time.
Next, for our LLMs we use all default parameters. For our LMMs, we experiment with modifying the maximum output token length as an ablation, but use all default parameters otherwise.
@@ B.1 Prompts
Our prompts are below. The black text is the base prompt template, and we replace the blue text with the corresponding information from the relevant video. The generated outputs are in the orange text.
Planner prompt $P_{T}$:
@@
Evaluator prompt $P_{V}$:
User: Evaluate if there is enough information to answer a multiple-choice question about a video and if the plan has been completed. If there is enough information to choose the correct answer with complete certainty and the plan has been followed, return the index of the choice after a brief explanation. Otherwise, return None after a brief explanation of why you can’t narrow down to a single answer choice. Be strict and don’t guess. INFORMATION: INFO PLAN: PLAN QUESTION: QUESTION CHOICES: CHOICES Give a brief explanation. Then, include your final answer after the words ”Final Answer:” in your response at the end. Do not include anything other than the answer as an integer or None after ”Final Answer:”. Let’s think step by step.   Assistant: Output: ANSWER

### v2

@@ 4.1 Datasets
We use the following datasets: (1) NExT-QA (Xiao et al., 2021) is a dataset that tests causal action reasoning and temporal understanding. Following the trend of works before us, we evaluate our method on the 5,000 questions in the NExT-QA validation set. (2) EgoSchema  (Mangalam et al., 2023) is a challenging dataset that tests long-form video understanding. Viewers need to view 100 seconds of the video on average to answer the question correctly. (3) STAR (Wu et al., 2021) tests reasoning in real-world video situations. (4) Perception Test (Patraucean et al., 2023) is a challenging dataset that focuses on skills such as memory, abstraction, physics, and semantics and is intended to be approached in a few-shot or zero-shot manner.
@@ 4.2 Baselines
In our experiments, we compare our method to recent state-of-the-art zero-shot (ZS) methods, such as LLoVi (Zhang et al., 2023a), ProViQ (Choudhury et al., 2023), and other methods that are not necessarily ZS, such as SeViLA (Yu et al., 2023), and MC-ViT (Balazevic et al., 2024). We note that SeViLA uses fine-tuned components on QV-Highlights (Lei et al., 2021a), while MC-ViT is fine-tuned on NExT-QA for Perception Test. Additional baselines are in Supplementary in Section A.2.
@@ 4.3 Results
Our results are shown in Table 1(d). We use GPT-4 for NExT-QA to ensure a fair comparison with LLoVi, which is the current state-of-the-art that uses GPT-4 to uniformly caption frames across the entire video. Interestingly, our method outperforms LLoVi despite viewing 50% fewer frames on average. Second, we also outperform SeViLA by +4.6%, although SeViLA uses a keyframe selector that is fine-tuned on a video moment retrieval and grounding task while our method is fully ZS.
For EgoSchema, we use GPT-4 with a retrained version of LaViLa that excludes overlapping Ego4D and EgoSchema videos to prevent data leakage. We show strong performance on long-form videos, where we outperform LLoVi by +3.0%, while viewing 95% fewer clips on average.
@@
We use GPT-3.5 for Perception Test and STAR because it is cheaper than GPT-4, but results are likely to be improved even further with GPT-4. Nevertheless, we achieve higher accuracy than LongViViT on Perception Test by +4.5% and MC-ViT by +2.1%, although it was fine-tuned on NExT-QA. On STAR, we surpass the best ZS approach by +2.7% and the best fine-tuned result by +0.3%.
@@
Finally, please refer to Section A for more experimental and ablation results, and Section D for additional visualizations. For example, in Figure 8, we see that question-answering is able to extract more relevant details in comparison to simple captioning. This may explain why our method significantly outperforms the descriptive split of NExT-QA. In Figure 7, we see our method is able to reason about vague references and correctly identify relevant objects through question-answering.
4.4 Agent Ablations
We perform ablations using 1000 randomly selected questions from the NExT-QA training set (see Table 2). Unless specified, we use GPT-3.5/LLaVA-1.6 as the LLM/LMM for all agents.
Ablating the Planner. The Planner module outputs a plan, a list of instructions that guides the behavior of all other modules. We test the impact of removing it from our framework and find that it is worse. We hypothesize this is because the Planner provides many temporal cues that guide the Retriever module’s search, such as “go to the middle of the video”, and without these cues, the Retriever is not as good at selecting the next timestamp. Moreover, the Planner also helps the Evaluator better decide when to stop since in our iterative approach, the Evaluator uses the plan to determine when to stop. We also try removing re-planning by only running the Planner once in the beginning and keeping this plan fixed throughout. We find that this also reduces performance, showing the importance of adjusting plans to new information.
Ablating the Retriever. The Retriever module determines the next timestamps to view, which helps focus our information collection. We ablate it by uniformly sampling frames from the video at 2-second intervals, similar to other methods like LLoVi, which performs worse. We believe this is because the Retriever allows us to capture frames that might have otherwise been skipped through uniform sampling and selects fewer unimportant frames that might mislead the model.
Table 2: Ablations on the agents in our framework.
Agent Ablation Avg. (%)
——– Baseline 60.4
Planner Removal 58.1
Fixed Plan 58.1
Retriever Removal 56.9
Extractor Caption Only 58.2
Evaluator No Iteration 56.8
Summarizer Removal 57.2
Ablating the Extractor. Question asking is important as it allows us to capture more fine-grained and question-relevant information compared to simple caption generation, which produces a generic description. Thus, we ablate the Extractor by only allowing the LMM to caption frames. We find that this decreases performance by -2.2%, suggesting that the ability to ask specific questions about a frame is important. We notice many generated captions capture the main idea of visual information in the frame, but are lacking in fine-grained details.
Ablating the Evaluator. The Evaluator reflects on the collected information and decides if there is enough information to answer the question. To examine the impact of this reflection process, we make the Evaluator answer on the first iteration, finding it results in a -3.6% performance decrease.
@@ 4.4 Agent Ablations
Table 3: Ablation Results. We perform ablations on 1000 randomly selected questions from the NExT-QA training set. We report (a), (b) replacing different LLMs and LMMs, (c) selecting different numbers of frames to view in the Retriever, and (d) changing the number of questions asked in the Extractor. We use GPT-3.5/LLaVA-1.6 as the LLM/LMM, 5 frames for the Retriever, and 3 questions for the Extractor, unless otherwise specified.
(a) Replacing Diff. LLMs
@@
4.5 Additional Experiments
@@
Substituting different LLMs/LMMs. To see how the choice of the LLM and LMM affects our framework’s performance, we swap different LLMs and LMMs into our framework (see Table 3(a) and Table 3(b)). We first try different LLMs while fixing the LMM to be LLaVA-1.6. While GPT-4 performs better than GPT-3.5 by a significant margin of +5.4%, open-source model Llama 3 is very close (-1.9%), while incurring no additional cost. We also evaluate the newly released GPT-4o, which outperforms GPT-4 by +2.2% while being 61% cheaper, showing that our method can leverage better future models. Second, we use different LMMs while fixing the LLM to be GPT-3.5. We find that LLaVA-1.6 does best, GPT-4V is slightly worse (-0.9%), and BLIP-2 is significantly worse (-7.7%). Finally, we run an experiment using GPT-4V as both the LLM and LMM, and find that this does worse than GPT-4 and LLaVA-1.6 by -1.1%.
GPT-4V Baseline. To get a baseline using GPT-4V, we use a method similar to LLoVi using a subset of 500 examples. We use GPT-4V to caption the video uniformly, then ask it to answer the question given the choices, which results in a performance change of -2.0%. We note our motivation is to refrain from captioning every single frame, instead finding frames that help us answer the question.
@@
Retriever window size. We experiment with different window sizes in Table 3(c) which is the number of frames the Retriever extracts centered around the selected frame. This allows us to capture better actions that occur quickly or require more context to understand. We find that choosing 5 frames yields the best results and a +1.4% increase when compared to selecting a single frame, but viewing more than 5 decreases performance. This suggests retrieving multiple frames can help the model better capture relevant information, but retrieving too many frames can lead to too much information.
Method NExT-QA (Random Subset)
Temporal Causal Descriptive Average
SeViLA - Localizer (Yu et al., 2023) 48.8 61.2 68.3 58.2
Moment-DETR (Lei et al., 2021a) 45.3 55.8 70.8 54.6
SigLIP (Zhai et al., 2023) 48.4 61.5 73.8 59.1
TraveLER - Planner & Retriever (ours) 50.9 62.7 72.4 60.3
Memory Intialization. In order for the Planner to create effective plans, it is beneficial to initialize the memory bank properly. Memory initialization allows the Planner to have a high-level overview of the video, and create a corresponding plan on how to traverse the video given the initial frames. We perform three different initializations with 1, 3, 5, 7, and 9 frames and display our results in  Figure 3. We observe that initializing the memory bank with 5 frames uniformly sampled from the video (0, 0.25, 0.5, 0.75, 1 for beginning, quarter, middle, three-quarter, end) yields the best result. In contrast, we notice a decrease in accuracy of -2.9% when using 9 frames, -3.1% when using 7 frames, -0.6% when initializing with 3 frames (0, 0.5, 1 for beginning, middle, end), and -2.5% when initializing with 1 frame (0.5 for middle).
Robustness to choosing incorrect frames. To see if our approach can recover from viewing incorrect frames, we choose random frames for the first 3 iterations (out of 5 total iterations) before using the Retriever. This leads to an accuracy drop of only 2.8%, showing our method is capable of recovering from viewing incorrect frames.
@@
The number of questions. Question answering allows us to extract more specific details from our visual inputs. However, we noticed that too many questions can yield irrelevant questions and false positives. Thus, we experiment with modifying the number of questions asked for each frame by our Extractor (see Table 3(d)). We record results for a 5-question, 3-question, and 1-question maximum. Note that 0-questions asked is equivalent to only allowing captions, which is discussed in the Extractor ablation. From our results, we find that a 3-question limit yields the best results compared to asking 1 or 5 questions (+2.0/+0.8%). This suggests that asking questions helps in extracting relevant information, but too many questions can lead to false positives or too much irrelevant information.
@@
Memory bank initialization and formatting. Our memory bank $M$ stores information that all modules rely on to make decisions. First, we experiment with different initializations as $M$ must be initialized in the first iteration. We experiment with initialization of 1, 3, and 5 uniformly sampled captions. We find using 5 evenly spaced frames yields the best results, possibly because it starts the model with a general overview of the video before it starts to collect more relevant information. We also experiment with changing the format of the memory bank from JSON to markdown tables. We find this does worse by -2.9%, possibly because LLMs are better able to understand JSON formats.
Cost of inference. We test costs for GPT-3.5 and GPT-4, which have costs per inference of \$0.03 and \$0.67, respectively. We also test the open-source model Llama 3, which is free to run with only a minor performance decrease (-1.9%). Finally, we also try the new GPT-4o model, which is both better (+2.2%) and cheaper (61%) than GPT-4. We believe this trend will only continue as models become better and cheaper in the future.
5 Conclusion
We design a modular, multi-LMM agent framework for video-question answering based on several agents with different roles, instructed by a Planner agent that updates its instructions using shared feedback between the other agents. Our method creates a plan to “traverse” through the video, asking questions about individual frames to “locate” and store key information, and then “evaluate” if there is enough information to answer the question, “replanning” using new feedback if necessary. Through extensive experiments and ablations, we find that the proposed TraveLER approach is not only easy to employ with different models but also improves performance on several video question-answering benchmarks without the need to fine-tune on specific datasets.
6 Limitations
In this work, we present a modular, zero-shot framework for video question answering (VideoQA) and demonstrate its effectiveness by improving on multiple state-of-the-art benchmarks. While TraveLER offers substantial benefits for VideoQA, it is important to recognize certain limitations that accompany our approach. Firstly, the effectiveness of our model relies heavily on the strength of the LLM and LMM. We notice that false-positives and incorrect statements from the LMM heavily impact performance. Our method’s runtime also depends on the runtime of current existing methods, and with the modularity of our method we expect this to improve with faster and better models in the future. Finally, we do not anticipate negative impacts of this work, but, as with any Machine Learning method, we recommend exercising caution.
7 Acknowledgements
@@
We would like to thank Suzie Petryk, Chancharik Mitra, Alon Mendelson, David Chan, Assaf Arbelle, and Leonid Karlinsky for helpful feedback and discussions. This project was supported in part by DoD, including PTG and/or LwLL programs, as well as BAIR’s industrial alliance programs.
@@ A.2 Additional Results
Figure 4: Comparison of different LMM Response Length (75, 150, 300 max tokens). 150 is optimal.
Results on Causal-VidQA. We also report zero-shot results on Causal-VidQA (Li et al., 2022), a dataset designed to focus on causal related questions to facilitate deeper video understanding towards video reasoning. We see that our method outperforms the best zero-shot method by 16.9%.
Additional Baselines. We also compare our approach with the concurrent work MoReVQA (Min et al., 2024). It can be seen that MoReVQA performs better on NExT-QA by 1%, while our method performs better on the more difficult long-form video understanding EgoSchema benchmark by 1.6%. We also note some key differences. Firstly, our method is able to re-plan based on feedback from previous iterations while MoReVQA uses a single forward pass through the stages and cannot modify its approach during the process. Secondly, we use an iterative approach while MoReVQA does not, which allows our method to adjust and improve using feedback from previous iterations.
Table 5: Zero-shot (ZS) results on Causal-VidQA.
Model Acc. (%)
Just-Ask (Yang et al., 2021) 27.1
CaKE + CoMem (Su et al., 2023) 30.0
CaKE + HGA (Su et al., 2023) 30.6
TraveLER (ours) 47.5 (+16.9)
@@ Appendix B Additional Implementation Details
To run our models on larger benchmarks, we use 8 NVIDIA RTX 6000 GPUs and split the dataset across multiple processes. In addition, we use the SGLang (Zheng et al., 2023) package, which provides a variety of performance optimizations for our LMMs and enables us to perform batched inference for models that do not natively support doing so. We serve our LMM on a single GPU and implement a queue that is shared across all runs. This allows individual runs to asynchronously call the LMM using an API request instead of creating a new instance of the LMM for each run. Typically, we have 4-5 processes sharing the same LMM.
Next, we use all default parameters. For our LMMs, we experiment with modifying the maximum output token length as an ablation, but use all default parameters otherwise. For the Llama 3 ablation in 3(a), we serve Llama 3 70B on 4 NVIDIA A100s using vLLM (Kwon et al., 2023). We report results from a single run for all experiments.
@@ B.1 Prompts
Our prompts are shown in Table 6, Table 7, Table 8, and 9. The black text is the base prompt template, and we replace the blue text with the corresponding information from the relevant video. The generated outputs are in the orange text.
@@
Table 8: Extractor prompt $P_{E}$:
User: Evaluate if there is enough information to answer a multiple-choice question about a video and if the plan has been completed. If there is enough information to choose the correct answer with complete certainty and the plan has been followed, return the index of the choice after a brief explanation. Otherwise, return None after a brief explanation of why you can’t narrow down to a single answer choice. Be strict and don’t guess. INFORMATION: INFO PLAN: PLAN QUESTION: QUESTION CHOICES: CHOICES Give a brief explanation. Then, include your final answer after the words "Final Answer:" in your response at the end. Do not include anything other than the answer as an integer or None after "Final Answer:". Let’s think step by step.   Assistant: Output: ANSWER
Table 9: Evaluator prompt $P_{V}$: